Methane (CH4) reacts with hydrogen sulfide to yield hydrogen gas and carbon disulfide, a solvent used in the manufacturing rayon and cellophane. What is the value of Kc at 1000 K if the partial pressures in an equilibrium mixture at 1000 K are 0.20 atm methane, 0.15 atm hydrogen sulfide, 0.30 atm carbon disulfide and 0.10 atm hydrogen gas? 


1
Expert's answer
2021-03-18T02:57:25-0400

CH4 + 2H2S = 4H+ CS2

Kp = "{(0.030)^4 (0.10) \over( 0.20)(0.15)^2} =1.8 \Chi 10^-5"


Kc = Kp/(RT)"^n" = "{1 .8 \Chi 10^-5\over (0.082 \Chi 1000)^2} = 2.68 \Chi 10^-9"